OpenStudy (anonymous):

If you are finding the line parallel to the graph of x + 3y = 5, what is the slope of the line?

OpenStudy (anonymous):

solve for y and theslope should be your answr which is 1/3

OpenStudy (aroub):

parallel=same slope 3y=-x+5 y=-1/3x+5/3 slope is -1/3

OpenStudy (asadkarim7):

parallel lines have same slope so 3y=-x+5 y = -1/3x+5 /3 the slope is -1/3

OpenStudy (anonymous):

y= /1/3 srry about the sign

OpenStudy (anonymous):

-1/3